# Frostman random variables, entropy inequalities, and applications

### v4

@@ 2.4 Frostman measures
(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$, if
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
@@
(ii) Write $p:=P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big)$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that, if $p>0$, then we must have
@@
One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P(X\in I_{i}^{\prime}),$
@@
$p$ $\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P\left(X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)\right)$
$\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}\cdot C_{1}\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)^{s_{1}}.$
@@
(i) The conclusion comes from [12, Lemma 4.8].
@@
The conclusion then follows.
@@ 2.5 Entropy of the distance function
for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [13, Section 4.2, page 61].
@@
$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O(\tilde{R}^{-t-1/2}):=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O(\eta^{t}).$
@@
In this subsection, we show some results of the discretized entropy of the distance function of certain measures, together with the square of the distance. These theorems are the key ingredients of the proof of Theorem 1.4.
@@
$H_{n}\big((X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big)\geq ns+O_{s}(\log_{+}C)+O_{s}(\log n),$
@@
$H_{n}\big(\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big)\geq ns+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{s}(\log n).$
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section.
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (24)
@@
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
@@
$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)+O(\log d)$
$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa)+O(\log d).$ (25)
@@
$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa)+O(\log d).$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
@@
$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)+O(\log d)$
$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (31)
@@
$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (35)
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$
@@
4 Proof of Theorems 1.21.4 with $s>1/2$
@@ 4.1 Diagonal quadratics with independent random variables
$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=O(\log\mathfrak{K}).\end{split}$ (36)
@@
$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})$-Frostman. Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\log\mathfrak{K})$ since $s_{1},s_{2}$ are bounded. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big(\sqrt{|\alpha|}c_{1}+\sqrt{|\beta|}c_{1}\big)$. So we can deduce by Lemma 2.32 and Theorem 2.41 that
$\begin{split}&H_{n}\big(X\big)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big(Y\big)\geq ns_{2}-\log C_{2}+O(1),\\ &\qquad\qquad\qquad H_{n}\big(|W_{1}-W_{2}|^{2}\big)\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (41)
for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows that
$n\left(s+\epsilon_{0}+5s_{1}+5s_{2}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
Recalling that $s_{1}+s_{2}\geq 2s$, we conclude that
@@
$\qquad\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (43)
Putting (42) and (43) together, we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
@@
$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))$
$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}).$
@@
$n\left(s+\epsilon_{0}+8s_{1}+8s_{2}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
Finally, the lemma will follow with constant $\epsilon_{1}=\epsilon_{0}/17$. ∎
@@
$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}x^{2}+\beta_{1}x+\beta_{0},$
@@
$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c).$
@@
Note that by completing the squares, we have
$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(x+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
@@
Let $\eta_{i}$ $(i=1,2)$ be parameters to be determined later, with $0<\eta_{i}\leq\min\{(4C_{i})^{-1/s_{i}},\,1/10\}$. Denote by $E$ the event that $|X+a|\geq\eta_{1}$, $|Y+b|\geq\eta_{2}$. Let $(X_{1},Y_{1})$ be the random variable such that $(X_{1},Y_{1})\equiv(X+a,Y+b\,|\,E)$ (see Lemma A.2). One can verify that $X_{1}$ and $Y_{1}$ are still independent. Since $(X+a,Y+b)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, one has
@@
$\begin{split}&H_{n}(X_{1}+Y_{1})+H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})\geq n(2s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log(1/\eta_{1}))+O(\log(1/\eta_{2}))+O(\log\mathfrak{K}).\end{split}$ (47)
@@
$\begin{split}&H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})\\ &\qquad\qquad\qquad+O\big(C_{1}\eta_{1}^{s_{1}}n+C_{2}\eta_{2}^{s_{2}}n+\log(1/\eta_{1})+\log(1/\eta_{2})\big).\end{split}$
Recall that we require $0<\eta_{i}<\min\{(4C_{i})^{-1/s_{i}},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$. Now assume that the implied constant in the last term on the right-hand side of above inequality is $C^{\prime}$, with $C^{\prime}\geq 1$. We then take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{4C^{\prime}C_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$ $(i=1,2)$. Then $C^{\prime}C_{i}\eta_{i}^{s_{i}}n\leq n\epsilon_{1}/4$, and $O(\log_{+}(1/\eta_{i}))=O(\frac{1}{s_{i}}\log_{+}C_{i})+O_{s}(1)$. Recall $s_{1},s_{2}\gg s$. Therefore, the proof is completed by taking $\epsilon_{1}^{\prime}=\epsilon_{1}/2$. ∎

### v5

@@ 2.4 Frostman measures
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
When $X$ is conditionally $(s,C)$-Frostman given $Y$, the conditional distribution of $X$ given $Y=y$ is $(s,C)$-Frostman for $P$-a.e. $y$.
@@
(ii) Write $p:=P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big)$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that if $p>0$, then we must have
@@
One can check that if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P(X\in I_{i}^{\prime}),$
@@
$p$ $\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P\left(X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)\right)$
$\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}\cdot C_{1}\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)^{s_{1}}.$
@@
(i) The conclusion comes from [19, Lemma 4.8].
@@
The conclusion then follows. ∎
@@ 2.5 Entropy of the distance function
for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [20, Section 4.2, page 61].
@@
$\qquad\ll\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}+\tilde{R}^{-t-1/2}:=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}+\eta^{t}\right).$
@@
In this subsection, we show some results of the discretized entropy of the distance function of certain measures, together with the square of the distance. These theorems are the key ingredients of the proof of Theorems 1.2 and 1.3.
@@
$H_{n}\big((X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big)\geq ns+O\left(\frac{1}{s}(\log_{+}C+\log n)+1\right),$
@@
$H_{n}\big(\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big)\geq ns+O\left(\frac{1}{s}(\log_{+}C+\log n)+1\right)+O(\log_{+}c).$
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [36, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem with weakly bi-Lipschitz functions.
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$ (25)
@@
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa+\log d).$
@@
$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa+\log d)$
$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa+\log d).$ (26)
@@
$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa+\log d).$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa+\log d).$
@@
$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa+\log d)$
$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa+\log d).$ (32)
@@
$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$ (36)
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$
@@
4 Proof of Theorems 1.2 and 1.3 with $s>1/2$
@@ 4.1 Diagonal quadratics with independent random variables
$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=H_{n}(X)+O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=H_{n}(Y)+O(\log\mathfrak{K}).\end{split}$ (37)
@@
$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$ (42)
Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})$-Frostman. Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\big|\log|\alpha|\big|+\big|\log|\beta|\big|)$ since $s_{1},s_{2}$ are bounded. Notice that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big((\sqrt{|\alpha|}+\sqrt{|\beta|})c_{1}\big)$. Also recall $s_{1}+s_{2}\geq 2s$. We can deduce by Lemma 2.32 and Theorem 2.41 that
$\begin{split}H_{n}(X)+H_{n}(Y)&\geq 2ns-\log C_{1}C_{2}+O(1),\\ H_{n}\big(|W_{1}-W_{2}|^{2}\big)&\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (43)
for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows from (42) and (43) that
$n\left(11s+\epsilon_{0}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
We conclude that
@@
$\qquad=\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (45)
Putting (44) and (45) together, we complete the proof of the lemma with $\epsilon_{1}=\epsilon_{0}/11$.
@@
$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$
$\qquad\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\mathfrak{K}).$
@@
$n\left(17s+\epsilon_{0}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
Finally, the lemma will follow with $\epsilon_{1}=\epsilon_{0}/17$. ∎
@@
$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}y^{2}+\beta_{1}y+\beta_{0},$
@@
$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c\big).$
Here, and from now on, the notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on other quantities such as $\rho_{1},\rho_{2},C_{1},C_{2},c$ or $n$.
@@
By completing the squares, we have
$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(y+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
@@
Let $\eta_{i}$ $(i=1,2)$ be parameters to be determined later, with $0<\eta_{i}\leq\min\big\{(4C_{i})^{-1/s_{i}},\,1/10\big\}$. Denote by $E$ the event that $|X+a|\geq\eta_{1}$, $|Y+b|\geq\eta_{2}$. Let $(X_{1},Y_{1})$ be the random variable such that $(X_{1},Y_{1})\equiv(X+a,Y+b\,|\,E)$ (see Lemma A.2). One can verify that $X_{1}$ and $Y_{1}$ are still independent. Since $(X+a,Y+b)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, one has
@@
$\begin{split}&H_{n}(X_{1}+Y_{1})+H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})\geq n(2s+\epsilon_{1})+O\big(\log(1/\eta_{1})+\log(1/\eta_{2})+\log\mathfrak{K}\big).\end{split}$ (49)
@@
$\begin{split}&H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})\\ &\qquad\qquad\qquad\qquad+O\big(C_{1}\eta_{1}^{s_{1}}n+C_{2}\eta_{2}^{s_{2}}n+\log(1/\eta_{1})+\log(1/\eta_{2})\big).\end{split}$ (50)
Recall that we require $0<\eta_{i}<\min\big\{(4C_{i})^{-1/s_{i}},1/10\big\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$. Now assume that the implied constant in the last term on the right-hand side of (50) is $C^{\prime}$, with $C^{\prime}\geq 1$. We then take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{4C^{\prime}C_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$ $(i=1,2)$. Then $C^{\prime}C_{i}\eta_{i}^{s_{i}}n\leq n\epsilon_{1}/4$, and $O(\log_{+}(1/\eta_{i}))=O\big(\frac{1}{s_{i}}\log_{+}C_{i}\big)+O_{s}\big(\frac{1}{s_{i}}\big)$. Recall $s_{1},s_{2}\gg s$. Therefore, the proof is completed by taking $\epsilon_{1}^{\prime}=\epsilon_{1}/2$. ∎